Circular No. 8715

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 K3 (McNAUGHT) R. H. McNaught reports his discovery of yet another comet on Uppsala

Schmidt telescope images (discovery observation tabulated below); the new object appeared diffuse with a 11′′ coma, extended 20′′ to the southeast; deeper exposures on May 22.37 UT show a well-condensed head with a 20′′ tail in p.a. 135^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
May 22.34471  6^{h}50^{m}29.^{s}50  −63^{o}44′05′′.7  17.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-K40.

T = 2007 Mar. 13.340 TT  ω = 327.^{o}144
Ω = 49.694 }2000.0
q = 2.54087 AU  i = 92.714

COMET C/2006 K4 (NEAT) An apparently asteroidal object reported by the NEAT project from

their CCD images taken with the 1.2-m Schmidt telescope at Palomar (discovery observation tabulated below) has been found cometary (following posting on the ‘NEO Confirmation Page’) by J. Young at Table Mountain (0.61-m f/16 Cassegrain reflector), who noted a round 5′′ coma with no central condensation and no tail on his images from May 24.4 UT.

2006 UT  α_{2000}  δ_{2000}  Mag.
May 18.45049  20^{h}53^{m}42.^{s}84  +14^{o}41′24′′.0  19.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-K41.

T = 2007 Dec. 3.052 TT  ω = 233.^{o}246
Ω = 116.524 }2000.0
q = 3.21880 AU  i = 111.517

COMET 73P/SCHWASSMANN-WACHMANN Further to IAUC 8709, additional components have been designated:

‘BQ’ on MPEC 2006-J31, and ‘BR’ and ‘BS’ on MPEC 2006-K18.

2006 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green